# SeNMo: A Self-Normalizing Deep Learning Model for Enhanced Multi-Omics Data Analysis in Oncology

### v1

@@ 3 Results
3.1 Pan-Cancer Multimodal Analysis - Overall Survival
@@
For the overall survival (OS) task, the pan-cancer data was randomly divided into the training-validation set ($80\%$) and the held-out test set ($20\%$), for each cancer type. The pan-cancer training was carried out by combining the training-validation cohort of all $33$ cancer types and adopting the 10-fold cross-validation with the $80-20\%$ division of samples. The training-validation cohort has $11,050$ patients, each having $\mathbb{R}^{80,697}$ features, comprised of the six multi-omics modalities including, gene expression, DNA methylation, miRNA expression, protein expression, DNA mutation, and the four clinical features (age, gender, race, stage). The SeNMo encoder model was trained on the training-validation cohort for the regression task of predicting the OS. C-Index was used as the evaluation metric of the hazard score predicted by the model. We used weights and biases [108] to find the optimal set of hyperparameters for our deep learning model. Figure 9 shows the visualization of the parallel sweeps across all hyperparameters, resulting in training around $400$ unique models. The best model had a learning rate of $0.00058$, a weight decay of $0.00598$, $0.1058$ dropout, $256$ batch size, $100$ epochs, and seven hidden layers with neurons in these layers as $[1024,512,256,128,48,48,48]$. The trained model has $83.33$ million trainable parameters. The checkpoints were saved for the best model for each of the $10$ folds. The model’s training resulted in the average training C-Index of $0.78$ and average validation C-Index of $0.76$, across the $10$ folds. For the evaluation of the trained model, the inference data was created by combining the held-out test set from all $33$ cancer types, resulting in $2754$ patients, each having $\mathbb{R}^{80,697}$ features. The inference on the test set showed the C-Index of $0.757$ which was the average of the C-Indices from the $10$ checkpoints. To further validate our findings, we created the ensemble of the $10$ checkpoints by averaging the prediction vectors from all the models and then evaluating the final averaged prediction vector for C-Index. For the pan-cancer, multi-omics SeNMo model, we got the ensemble C-Index of $0.758$ on the held-out test set. The significance level in all these analyses is $95\%$, i.e., $p<0.05$ which indicates statistically significant values. These results are depicted in the Figure 5. We further tested the optimal hyperparameters of our trained model to train different combinations of the pan-cancer data modalities. We call these 1-modal, 3-modal (gene expression, DNA methylation, miRNA expression), 4-modal ($3+$protein expression), 5-modal ($4+$DNA mutation), and 6-modal (all modalities) cohorts. It is important to note that our initially trained pan-cancer, multi-omics model belongs to the 6-modal cohort. The purpose here is to see how the model performs on each of these pan-cancer cohorts where one or more of the data modalities is missing. As depicted in Figure 5, the SeNMo model trained on the pan-cancer 1-modal (Gene expression) cohort showed a C-Index for training, validation, testing, and ensemble inference as $0.729,0.702,0.718,$ and $0.728$, respectively. For the pan-cancer 1-modal (DNA methylation) cohort, the model’s training, validation, testing, and ensemble inference C-indices are $0.636,0.629,0.644,$ and $0.65$, respectively. For the pan-cancer 1-modal (miRNA expression) cohort, the model’s training, validation, testing, and ensemble inference C-indices are $0.744,0.68,0.686,$ and $0.702$, respectively. We did not analyze the model individually on the rest of the three modalities because clinical and protein expression features are too small for an $83$ million-parameter model, whereas the DNA mutation data comprised the binarized features of mutations. Evaluating the model on the 3-modal cohort showed the training, validation, testing, and ensemble inference C-indices of $0.783,0.727,0.725,$ and $0.726$, respectively. Further adding the protein expression to the 3-modal data, we trained and evaluated the model on the 4-modal cohort and got the C-Indices of $0.88,0.742,0.746,$ and $0.751$ for training, validation, testing, and ensemble inference, respectively. Lastly, the model’s performance on the 5-modal cohort showed the training, validation, testing, and ensemble inference C-indices of $0.885,0.741,0.746,$ and $0.749$, respectively. Next, we analyze how the model trained on pan-cancer, 6-modal data fared on individual cancer patients’ data.
3.2 Cancer-wise Multimodal Analysis - Overall Survival
We evaluated the model trained on the 6-modal pan-cancer cohort on the held-out individual cancer data. The number of patients in these cancer cohorts was a randomly selected subset of the cases shown in Figure 3, and Tables 1, 2, which accounts for the $20\%$ of the total samples. The trained model was evaluated on each of the 33 individual cancer data using simple inference as well as the ensemble of the 10-fold checkpoints. Figure 6 shows the evaluation performance of the model on 33 cancer types. The model showed the best predictive performance on TCGA-PCPG data with an average C-Index on the test set of $0.9$ and ensemble inference of $0.929$. SeNMo’s performance on the e cancer types in format $\{TestInference,EnsembleInference\}$ is; TCGA-ACC=$=\{0.866,0.861\}$, TCGA-UVM=$=\{0.822,0.829\}$, TCGA-LGG=$=\{0.821,0.823\}$, TCGA-KICH=$=\{0.801,0.807\}$, TCGA-KIRC=$=\{0.777,0.776\}$, TCGA-KIRP=$=\{0.775,0.778\}$, TCGA-UCEC=$=\{0.708,0.713\}$, TCGA-THCA=$=\{0.696,0.698\}$, TCGA-SKCM=$=\{0.691,0.689\}$, TCGA-BRCA=$=\{0.687,0.692\}$, TCGA-CESC=$=\{0.676,0.682\}$, TCGA-ESCA=$=\{0.650,0.648\}$, TCGA-LUAD=$=\{0.647,0.653\}$, TCGA-SARC=$=\{0.65,0.658\}$, TCGA-STAD=$=\{0.631,0.628\}$, TCGA-LIHC=$=\{0.627,0.629\}$, TCGA-BLCA=$=\{0.609,0.609\}$, TCGA-MESO=$=\{0.599,0.615\}$, TCGA-LUSC=$=\{0.588,0.592\}$, TCGA-PAAD=$=\{0.597,0.598\}$, TCGA-HNSC=$=\{0.583,0.583\}$, TCGA-CHOL=$=\{0.574,0.574\}$, TCGA-COAD=$=\{0.546,0.542\}$, TCGA-THYM=$=\{0.555,0.571\}$, TCGA-UCS=$=\{0.514,0.541\}$, TCGA-OV=$=\{0.518,0.509\}$, TCGA-GBM=$=\{0.495,0.493\}$, TCGA-LAML=$=\{0.482,0.485\}$, TCGA-DLBC=$=\{0.714,0.619\}$, TCGA-READ=$=\{0.55,0.551\}$, TCGA-PRAD=$=\{0.304,0.3\}$, TCGA-TGCT=$=\{0.123,0.091\}$. We noticed that the results for TCGA-GBM, TCGA-LAML, TCGA-PRAD, and TCGA-TGCT were not statistically significant, i.e., $p>0.05$. So, we fine-tuned the model for these datasets by reducing the learning rate, increasing the weight decay and dropout, and letting the model fine-tune for 10 epochs. Resultantly, the model’s performance increased for TCGA-GBM=$=\{0.642,0.650\}$, TCGA-LAML=$=\{0.627,0.626\}$, and TCGA-PRAD=$=\{0.541,0.542\}$. These improvements are depicted with the green arrows and green boxes in Figure 6. However, the model failed to converge for TCGA-TGCT data and consistently gave insignificant predictions.
Figure 6: Cancer-wise C-Index results for OS prediction for some of the cancer types.
@@ 3.3 Out-of-distribution Evaluation and Fine-tuning
To further verify the performance of our model, we evaluated the model with the off-the-shelf datasets CPTAC-LSCC [56] and Moffitt’s LSCC [57]. These datasets were not part of the training data, depicting the out-of-distribution evaluation. Evaluating the model without fine-tuning showed the $\{TestInference,EnsembleInference\}$ of CPTAC-LSCC=$=\{0.48,0.50\}$, and Moffit-LSCC=$=\{0.581,0.59\}$. Fine-tuning the model for 10 epochs, with reduced learning rate, and increased weight decay and dropout resulted in the improvement of C-Indices as, CPTAC-LSCC=$=\{0.677,0.73\}$, and Moffit-LSCC=$=\{0.647,0.656\}$. These fine-tuning results are depicted in Figure 6 as the green box plots.
Figure 7: Kaplan-Meier Curves by Risk Group. KM comparative analysis of using SeNMo in stratifying patient outcomes in low/ intermediate/ high risk, defined by the 33-66-100 percentile of hazard predictions. Hazard predictions from SeNMo show clear distinctions between stratified groups. The p-values from logrank test for Low vs. Intermediate: $051.66e-$ 05, Low vs. High: $461.156e-$ 46, and Intermediate vs. High: $221.92e-$ 22. The shaded areas around each curve depicts the 95% confidence intervals.
@@ 3.4 Patient Stratification
Risk stratification of patients allows clinicians and researchers to identify patients who might need more intensive care or monitoring and those who may have a better prognosis, facilitating more personalized treatment approaches. We further investigate the SeNMo’s ability to stratify the patients based on low, intermediate, and high risk conditions. We plot Kaplan-Meier (KM) curves of our model on the pan-cancer, multi-omics held-out test set, as shown in Figure 7. We select the low/ intermediate/ high risk stratification distribution as the 33-66-100 percentile of hazard predictions [102, 109]. The hazard scores predicted by SeNMo are used to evaluate the model’s stratification ability. The KM comparative analysis shows that SeNMo distinguished the patients in the three groups across. The low-risk group (green) exhibited the highest survival probability, maintaining close to 100% survival up to approximately 5 years, and gradually declining to about 60% by the 25-year mark. The intermediate-risk group (blue) showed a significantly lower survival probability, starting to diverge from the low-risk group early on and reaching around 40% by the end of the study period. The high-risk group (orange) displayed the most pronounced decline in survival probability, with a steep drop to approximately 20% survival within the first 10 years, and further reducing to below 10% after 10 years. The logrank test to evaluate the significance of this stratification shows that the p-value of low vs. intermediate curves is $051.66e-$ 05, low vs. high is $461.156e-$ 46, and intermediate vs. high is $221.92e-$ 22, showing significant results, i.e., $p<0.05$. The 95% confidence intervals around each curve show the reliability of these estimates.
@@ 3.5 Primary Cancer Type Prediction
To test the generalizability of SeNMo across different tasks, we carried out the prediction of primary cancer type from pan-cancer, multi-omics data. We set the problem as a classification problem, where the multi-omics data is used to predict the type of cancer for the given patient data among the 33 classes. It is imperative to mention here that the four clinical features in the initial data contained the cancer stage, as shown in Figure 2 and Table 2. When considering cancer type classification problem, the stage adds a bias in the data because of the staging distribution among different cancers. So, for the cancer classification simulations, we excluded the “stage" feature in the clinical data. As shown in Figure 8, the model achieves near-perfect accuracy levels, with 99.9% average accuracy in training, 99.8% in validation, and consistent performance in both simple and ensemble inference approaches. The confusion matrix depicts a clear concentration of values along the diagonal, indicating a high rate of correct predictions across all cancer types. The scatter plot shows an alignment of predicted labels with true labels along the diagonal line, highlighting the model’s robust predictive accuracy. The classification report across various cancer types reveals that the model consistently maintains high precision, recall, and F1-scores, approaching a value of 1 for almost all categories.
@@
Figure 8: Pan-cancer primary cancer type prediction. The model’s accuracy across training, validation, and inference stages is near-perfect (top left panel). Confusion matrix (bottom left) shows minimal misclassifications, while the scatter plot (bottom middle) shows the alignment of predicted versus true labels. The classification report (right panel) shows high precision, recall, and f1-scores in the 33 cancers type-identification.
4 Discussion and Conclusion
@@
We analyzed pan-cancer dataset of 33 cancer types comprising five molecular data modalities having varying features and four features of clinical data using SeNMo encoder-based framework. In training such a large dataset having high-dimensional heterogeneous data is challenging. Molecular data, such as gene expression, miRNA expression, DNA methylation and mutations, and proteomics, are available from the same patient. This high-dimensional data has intra- and inter-dataset correlations, heterogeneous measurement scales, missing values, technical variations, and other forms of noise [10]. The optimal parameters of a deep learning model that can learn from this data is a non-trivial task. After extensive training-evaluation runs, we found a model that has shown to perform very well across the different data types and tasks, refer to Figures 5 and 9. The model has been shown to outperform the existing works in OS prediction when considering the six data modalities [36]. Moreover, we observe that adding more numbers and types of modalities increased the model’s performance.
@@
The model’s performance was evaluated on individual cancers at test-time through simple inference and ensembling methods. We observed that the model’s predictive power improved when an ensemble of the checkpoints was employed, refer to Figure 6. However, for four data types, TCGA-GBM, TCGA-LAML, TCGA-PRAD, and TCGA-TGCT, the model showed insignificant predictive power. During the investigation, we observed that these datasets had non-admissible pairs in some of the data folds, i.e., all samples had censor value $\delta=0$ in Equation 10. In the case of TCGA-PRAD and TCGA-TGCT, the number of samples having $\delta=1$ in the training/validation cohort was $12$ and $3$, respectively. We fine-tuned the model for these datasets by using the stratified k-folds to offset the class-representation problem in the data folds. After searching for the optimal hyperparameters for fine-tuning, the model’s performance became significant ($p<0.05$) for three out of four datasets, refer to green box plots in Figure 6. The model failed to converge for the TCGA-TGCT data despite numerous efforts at fine-tuning. The authors are of the view that because of the more pronounced class imbalance in this dataset, i.e., 3 out of 111 samples in the training cohort have censor value $\delta=1$, the model cannot learn to distinguish the samples based on the OS.
@@
It is imperative to mention here that MLPs-based networks are very sensitive to catastrophic forgetting when presented with out-of-distribution data or when subjected to a different task [110]. We fine-tuned the SeNMo encoder for one public data (CPTAC-LSCC) and one internal data (Moffitt’s LSCC) [56, 57]. In our simulations to fine-tune the model, we encountered the catastrophic forgetting phenomenon in SeNMo, where the model would fail to converge on these datasets. This was more pronounced when a certain number of hidden layers were frozen, and the rest were trained with lower learning rates, as shown in Figures 10 and 11. We resorted to the option of unfreezing all the layers of the encoder and fine-tuning the model with a very small learning rate ($54e-$ 5), high weight decay and dropout ($0.35$), and just $10$ epochs. This method worked and the model showed significant performance on the out-of-distribution datasets.
@@
The KM survival curves depicted in Figure 7 demonstrate a clear stratification of survival probabilities among three risk-defined patient groups. These results underscore the effectiveness of the risk stratification model in predicting long-term outcomes and highlight the critical need for targeted therapeutic strategies based on individual risk assessments. This stratification allows for more personalized patient management and could potentially guide clinical decision-making toward improving OS rates across diverse patient populations.
@@
The classification results in Figure 8 illustrate the superior generalizability to the prediction of primary cancer types of the SeNMo encoder, which was primarily trained for predicting OS. Furthermore, the detailed classification report across various cancer types reveals that the model consistently maintains high precision, recall, and F1-scores, for almost all categories. Such metrics not only confirm the model’s effectiveness in accurately identifying the correct cancer class but also its reliability in replicating these results across different samples. This level of performance suggests the capability of the model to handle complex patterns in data.

### v2

@@ 3 Results
3.1 Pan-Cancer Multimodal Analysis for Predicting Overall Survival
@@
Figure 11 shows the visualization of the parallel sweeps across all hyperparameters, resulting in training around $400$ unique models. The optimal model had a learning rate of $0.00058$, a weight decay of $0.00598$, $0.1058$ dropout, $256$ batch size, $100$ epochs, and seven hidden layers with neurons in these layers as $[1024,512,256,128,48,48,48]$. The trained model contained $83.33$ million trainable parameters. Checkpoints were saved for this model for each of the $10$ folds. The model’s training resulted in the average training C-Index of $0.78$ and average validation C-Index of $0.76$ across the $10$ folds. The inference on the test set showed the C-Index of $0.757$, the average of the C-Indices from the $10$ checkpoints. To further validate our findings, we created an ensemble of the $10$ checkpoints by averaging the prediction vectors from all the models and then evaluating the final averaged prediction vector for C-Index. For the pan-cancer, multi-omics SeNMo model, an ensemble C-Index of $0.758$ was achieved on the held-out test set. The significance level in all these analyses is $95\%$, i.e., $p<0.05$, indicating statistically significant values. These results are depicted in the Figure 6.
As depicted in Figure 6, the SeNMo model trained on the pan-cancer 1-modal (Gene expression) cohort showed a C-Index for training, validation, testing, and ensemble inference as $0.729,0.702,0.718,$ and $0.728$, respectively. For the pan-cancer 1-modal (DNA methylation) cohort, the model’s training, validation, testing, and ensemble inference C-indices are $0.636,0.629,0.644,$ and $0.65$, respectively. For the pan-cancer 1-modal (miRNA expression) cohort, the model’s training, validation, testing, and ensemble inference C-indices are $0.744,0.68,0.686,$ and $0.702$, respectively. We did not analyze the model individually on the rest of the three modalities because clinical and protein expression features are too small for an $83$ million-parameter model, whereas the DNA mutation data comprised the binarized features of mutations. Evaluating the model on the 3-modal cohort showed the training, validation, testing, and ensemble inference C-indices of $0.783,0.727,0.725,$ and $0.726$, respectively. Further adding the protein expression to the 3-modal data, we trained and evaluated the model on the 4-modal cohort and got the C-Indices of $0.88,0.742,0.746,$ and $0.751$ for training, validation, testing, and ensemble inference, respectively. Lastly, the model’s performance on the 5-modal cohort showed the training, validation, testing, and ensemble inference C-indices of $0.885,0.741,0.746,$ and $0.749$, respectively. Next, we analyze how the model trained on pan-cancer, 6-modal data fared on individual cancer patients’ data.
Figure 7: Cancer-specific C-index results for overall survival (OS) prediction across multiple cancer types. Pink box plots represent predictions on the held-out test set, while green box plots show predictions from fine-tuned models for cancer types with initially insignificant results or for unseen data, such as Moffitt LSCC and CPTAC-LSCC. Triangular markers indicate ensemble predictions, with blue triangles representing fine-tuned ensemble results. Models with insignificant predictions fall below the 0.5 threshold, marked by a red arrow. Although trained on pan-cancer cohort, SeNMo effectively captures survival times across individual cancers, with fine-tuning improving performance for cases with initially low predictive significance.
3.2 Individual Cancer Multimodal Analysis for Predicting Overall Survival
[…]
Table 5: C-Index for Test and Ensemble Inference across Cancer Types.
Cancer Type C-Index {Test, Ensemble} Cancer Type C-Index {Test, Ensemble}
TCGA-PCPG {0.900, 0.929} TCGA-BLCA {0.609, 0.609}
TCGA-ACC {0.866, 0.861} TCGA-MESO {0.599, 0.615}
TCGA-UVM {0.822, 0.829} TCGA-LUSC {0.588, 0.592}
TCGA-LGG {0.821, 0.823} TCGA-PAAD {0.597, 0.598}
TCGA-KICH {0.801, 0.807} TCGA-HNSC {0.583, 0.583}
TCGA-KIRC {0.777, 0.776} TCGA-CHOL {0.574, 0.574}
TCGA-KIRP {0.775, 0.778} TCGA-COAD {0.546, 0.542}
TCGA-UCEC {0.708, 0.713} TCGA-THYM {0.555, 0.571}
TCGA-THCA {0.696, 0.698} TCGA-UCS {0.514, 0.541}
TCGA-SKCM {0.691, 0.689} TCGA-OV {0.518, 0.509}
TCGA-BRCA {0.687, 0.692} TCGA-GBM {0.495, 0.493}
TCGA-CESC {0.676, 0.682} TCGA-LAML {0.482, 0.485}
TCGA-ESCA {0.650, 0.648} TCGA-DLBC {0.714, 0.619}
TCGA-LUAD {0.647, 0.653} TCGA-READ {0.550, 0.551}
TCGA-SARC {0.650, 0.658} TCGA-PRAD {0.304, 0.300}
TCGA-STAD {0.631, 0.628} TCGA-TGCT {0.123, 0.091}
TCGA-LIHC {0.627, 0.629}
@@ 3.3 Out-of-distribution Evaluation and Fine-tuning
[…]
Figure 8: Kaplan-Meier (KM) comparative analysis of using SeNMo in stratifying patient outcomes in low/ intermediate/ high risk, defined by the 33-66-100 percentile of hazard predictions. Hazard predictions from SeNMo show clear distinctions between stratified groups. The p-values from logrank test for Low vs. Intermediate: $051.66e-$ 05, Low vs. High: $461.156e-$ 46, and Intermediate vs. High: $221.92e-$ 22. The shaded areas around each curve depicts the 95% confidence intervals.
@@ 3.4 Patient Stratification
We further investigated the SeNMo’s ability to stratify the patients based on low, intermediate, and high risk conditions. We generate Kaplan-Meier (KM) curves of our model on the pan-cancer, multi-omics held-out test set, as shown in Figure 8. We select the low/ intermediate/ high risk stratification distribution as the 33-66-100 percentile of hazard predictions [101, 111]. The hazard scores predicted by SeNMo are used to evaluate the model’s stratification ability. The KM comparative analysis shows that SeNMo distinguished the patients across the three groups. The low-risk group (green) exhibited the highest survival probability, maintaining close to 100% survival up to approximately 5 years, and gradually declining to about 60% by the 25-year mark. The intermediate-risk group (blue) showed a significantly lower survival probability, starting to diverge from the low-risk group early on and reaching around 40% by the 15-year mark of the study period. The high-risk group (orange) displayed the most pronounced decline in survival probability, with a steep drop to approximately 20% survival within the first 10 years, and further reducing to below 10% after 10 years. The logrank test to evaluate the significance of this stratification shows that the p-value of low vs. intermediate curves is $051.66e-$ 05, low vs. high is $461.156e-$ 46, and intermediate vs. high is $221.92e-$ 22, showing significant results, i.e., $p<0.05$. The 95% confidence intervals around each curve show the reliability of these estimates.
@@ 3.5 Primary Cancer Type Prediction
To test the generalizability of SeNMo across different tasks, we carried out the prediction of primary cancer type from pan-cancer, multi-omics data. We set the problem as a classification problem, where the multi-omics data is used to predict the type of cancer for the given patient data among the 33 classes. It is imperative to mention here that the four clinical features in the initial data contained the cancer stage, as shown in Figure 2 and Table 2. When considering a cancer type classification problem, the stage adds a bias in the data because of the staging distribution among different cancers. Therefore, for the cancer classification simulations, we excluded the “stage” feature in the clinical data. As shown in Figure 9, the model achieves near-perfect accuracy levels, with 99.9% average accuracy in training, 99.8% in validation, and consistent performance in both simple and ensemble inference approaches. The confusion matrix depicts a clear concentration of values along the diagonal, indicating a high rate of correct predictions across all cancer types. The scatter plot shows an alignment of predicted labels with true labels along the diagonal line, highlighting the model’s robust predictive accuracy. The classification report across various cancer types reveals that the model consistently maintains high precision, recall, and F1-scores, approaching a value of 1 for almost all categories. The robust predictive power of our model emphasizes the fact that each cancer has a unique molecular landscape, highlighted through differences in gene, protein, and miRNA expression, DNA methylation, and types of somatic mutations seen in our data.
@@
Figure 9: Pan-cancer primary cancer type prediction results. The model’s accuracy across training, validation, and inference stages is near-perfect (top left panel). Confusion matrix (bottom left) shows minimal misclassifications, while the scatter plot (bottom middle) shows the alignment of predicted versus true labels. The classification report (right panel) shows high precision, recall, and f1-scores in the 33 cancers type-identification.
Figure 10: Tertiary Lymph Structures (TLS) Ratio Predictions by SeNMo. (a) Dataset distribution, dividing 103 samples into training, validation, test, and excluding those missing TLS data. (b) Box plots compare TLS ratios between manual annotations (M) and SeNMo predictions, showing no significant difference ($p=0.1$). (c) Violin plot of TLS Ratio annotations for high vs. low groups thresholded by the median value, with a significant difference ($p<0.05$). (d) Violin plot of SeNMo’s TLS Ratio predictions, also showing significant separation between high and low groups ($p<0.05$). (e) Kaplan-Meier survival curves with significant survival differences, comparing high vs. low TLS ratios for both annotations (top, $p=0.019$) and SeNMo predictions (bottom, $p=2.5e-4$).
3.6 Tertiary Lymph Structures (TLS) Ratio
To further evaluate SeNMo’s generalizability on previously unseen data and across different tasks, we fine-tuned the model to predict the TLS ratio on a cohort of lung squamous cell carcinoma data collected at Moffitt Cancer Center. This task was formulated as a regression problem. As shown in Figure 10, the TLS ratio predictions generated by SeNMo demonstrated strong performance on the held-out test set. Specifically, the comparison between manual TLS ratio annotations and SeNMo-predicted ratios revealed no significant difference ($p=0.1$), indicating a high level of concordance between manual assessments and model predictions (Figure 10b). Further analysis using violin plots compared the distribution of TLS ratios for manually annotated high vs. low groups with those predicted by SeNMo. Both manual and predicted TLS ratios showed significant separation between high and low groups ($p<0.05$), highlighting the model’s ability to accurately distinguish between different levels of TLS (Figures 10c and 10d). Moreover, KM survival analysis was performed to assess the prognostic value of TLS ratios. Survival curves revealed significant differences in survival outcomes between patients with high and low TLS ratios, both for manually annotated data ($p=0.019$) and SeNMo-predicted data ($p=2.5e-4$) (Figure 10e).
4 Discussion
@@
We analyzed pan-cancer dataset of 33 cancer types comprising five molecular data modalities (with varying amount of features) and four clinical data features using our SeNMo encoder-based framework. Public databases such as CPTAC and TCGA contain common identifiers within their data that connect data from the same patient. Therefore, molecular data, such as gene expression, miRNA expression, DNA methylation, somatic mutations, and protein expression can be consolidated to represent a singular patient. However, such high-dimensional data has intra- and inter-dataset correlations, heterogeneous measurement scales, missing values, technical variations, and other forms of noise [10]. This necessitates the need for a variety of preprocessing techniques such as the removal of low variance features and the imputing of missing features among others prior to training. Training such a large dataset having high-dimensional heterogeneous data required proper computational resources and a precise pipeline for training, testing, and validation. After extensive training-evaluation runs, we found, through optimal parameters searching, a model that performs very well across the different data types and tasks (refer to Figures 6 and 11). The model has been shown to outperform the existing works in OS prediction when considering the six data modalities included in our data [37]. Moreover, we observed that adding more data and types of modalities increased the model’s performance.
@@
The model’s performance was evaluated on individual cancers at test-time through simple inference and ensembling methods. We observed that the model’s predictive power improved when an ensemble of the checkpoints was employed, (refer to Figure 7). However, for four cancer types, TCGA-GBM, TCGA-LAML, TCGA-PRAD, and TCGA-TGCT, the model did not show significant predictive power. During the investigation, we observed that these datasets had non-admissible pairs in some of the data folds, i.e., all samples had censor value $\delta=0$ in Equation 10. In the case of TCGA-PRAD and TCGA-TGCT, the number of samples having $\delta=1$ in the training/validation cohort was $12$ and $3$, respectively. To address the lack of predictive power, we fine-tuned the model for these datasets by using the stratified k-folds to offset the class-representation problem in the data folds. After searching for the optimal hyperparameters for fine-tuning, the model’s performance became significant ($p<0.05$) for three out of four datasets, (refer to green box plots in Figure 7).
@@
It is imperative to mention here that MLPs-based networks are very sensitive to catastrophic forgetting when presented with out-of-distribution data or when subjected to a different task [112]. We fine-tuned the SeNMo encoder for one public data (CPTAC-LSCC) and one internal data (Moffitt’s LSCC) [57, 58]. In our simulations to fine-tune the model, we encountered the catastrophic forgetting phenomenon in SeNMo, where the model would fail to converge on both new datasets. This was more pronounced when a certain number of hidden layers were frozen, and the rest were trained with lower learning rates. We resorted to the option of unfreezing all the layers of the encoder and fine-tuning the model with a very small learning rate ($54e-$ 5), high weight decay and dropout ($0.35$), and just $10$ epochs. This method worked and the model showed significant performance on the out-of-distribution datasets.
@@
Risk stratification of patients allows clinicians and researchers to identify patients who might need more intensive care or monitoring and those who may have a better prognosis, facilitating more personalized treatment approaches. The KM survival curves depicted in Figure 8 demonstrate a clear stratification of survival probabilities among three risk-defined patient groups. These results underscore the effectiveness of the risk stratification model in predicting long-term outcomes and highlight the critical need for targeted therapeutic strategies based on individual risk assessments. This stratification allows for more personalized patient management and could potentially guide clinical decision-making toward improving OS rates across diverse patient populations.
@@
Cancer type classification is routinely studied for early detection and localization of tissue of origin [113]. The classification results in Figure 9 illustrate the superior generalizability of the model’s predictive power to classify primary cancer types through the SeNMo encoder, despite it being primarily trained for predicting OS. Additionally, the detailed classification report across various cancer types reveals that the model consistently maintains high precision, recall, and F1-scores for almost all cancer types. Such metrics not only confirm the model’s effectiveness in accurately identifying the correct cancer class but also its reliability in replicating these results across different samples. This level of performance suggests the capability of the model to successfully learn high level representations from heterogenous, high-dimension, mutlivariate data stemming from complex molecular modalities such as gene expression, miRNA expression, somatic mutations, DNA methylation, and protein expression.
As shown in Figure 10, SeNMo’s ability to predict TLS ratios was evaluated on an unseen cohort of lung squamous cell carcinoma data from Moffitt Cancer Center. The comparison between manual TLS ratio annotations and SeNMo-predicted values showed no significant difference ($p=0.1$), indicating a high level of concordance between human annotations and model predictions. Violin plots depicting high vs. low TLS ratio groups—both for manual and SeNMo predictions—revealed significant separation ($p<0.05$), demonstrating the model’s robustness in distinguishing between biologically distinct TLS levels. Furthermore, KM survival curves for high vs. low TLS ratio groups revealed significant differences in survival outcomes, with stronger statistical significance observed for SeNMo-predicted data ($p=2.5e-4$) compared to manual annotations ($p=0.019$). These results underscore the potential of SeNMo to not only replicate expert-driven TLS annotations but also provide a consistent and potentially superior prognostic assessment. Overall, the results indicate that SeNMo can successfully generalize to new tasks and datasets, accurately predicting TLS ratios and offering valuable prognostic insights that could improve clinical decision-making.
We made the entire codebase of SeNMo publicly available on GitHub (https://github.com/lab-rasool/SeNMo). We have made the latent representations of patient data generated from SeNMo available to the research community through our HoneyBee system [114]. HoneyBee stores these representations, also known as patient embeddings, in a structured format using Hugging Face datasets, effectively creating a vector database. HoneyBee has demonstrated the effectiveness of using patient embeddings, offering a significant advantage over the traditional approach of using raw data and extensive pre-processing [114].
5 Conclusion
In this study, we introduced SeNMo, a foundational deep learning model specifically designed for multi-omics data analysis across 33 different cancer sites. By leveraging high-dimensional multi-omics datasets from the NCI Genomics Data Commons, SeNMo demonstrated robust performance in predicting overall survival on both training and held-out test sets. The model’s adaptability and efficiency were further validated through its high accuracy in classifying primary cancer types and predicting TLS ratios, showcasing its ability to generalize effectively across different tasks. As a foundational model, SeNMo represents a resilient and scalable solution that advances the integration and analysis of complex molecular data, providing a comprehensive understanding of cancer biology. Our approach underscores the potential of self-normalizing networks in oncology, emphasizing the importance of comprehensive data preprocessing and optimal parameter tuning. By making SeNMo and its derived patient embeddings publicly available, we aim to facilitate further research and innovation in personalized cancer care, underscoring the transformative potential of multi-omics approaches in the fight against cancer.